OpenStudy (anonymous):

write and exponential function y=ab^x for a graph that includes (0,6) and (1,15)

OpenStudy (anonymous):

well anything to the zeroth power is 1 so you know a = 6. then you need y=6*b^x to have the point (1,15) so plug in your known x and y values and solve for b. 15 = 6*b^1 b^1 = b = 15/6 = 5/2 = 2.5 y = 6 * 2.5^x

OpenStudy (anonymous):

thank you so much! @LifeEngineer